What is the accumulated value after 15 years of monthly deposits of $ 1500.00 earning interest at 8% compounded monthly if the deposits are made at the beginning of each month?

Select one:

a. $513 102.31

b. $522 536.41

c. $518 507.23

d. $516 467.23

e. $519 075.73
Answer #1

b. $522 536.41

Using a financial calculator

Set the mode to BGN mode => 2nd BGN => 2nd SET => 2nd QUIT

PV = 0

PMT = 1500

N = 180 (15 years*12 contributions per year = 180 periods)

I/Y = 8/12 (8%/12 rate per monthly period)

cpt FV, we get FV = 522536.41

Hence, after 15 years $522536.41is accumulated